# Quantifying Energy and Cost Benefits of Hybrid Edge Cloud: Analysis of Traditional and Agentic Workloads

### v1

@@ 4.1 TRADITIONAL WORKLOADS (BASELINE)
Table 1 and Figure 1 show the costs and energy benefits of HEC compared with traditional cloud-based solutions.
![[Uncaptioned image]](extracted/6136269/Table_1.png)
@@
Without HEC, the centralized cloud energy consumption is about: 1,927 kWh/device/year. With HEC and assuming an 80 edge split, the energy consumption is approximately 674 kWh/device/year. The resulting savings with HEC is approximately: 62.
@@
With HEC and assuming an 80 edge split, the total costs of hosting, bandwidth and software licenses is about 66 per device per year. The resulting savings with HEC is about 200 per device per year or approximately 75.
For agentic workloads the percentage savings is the same but the savings per device per year are almost an order of magnitude larger providing up to 10,000 kWh per year in energy and 1,500 in bandwidth and hosting cost per year per device which emphasizes why HEC is indispensable to help make agentic economy sustainable.

### v2

@@ 4.1 TRADITIONAL WORKLOADS (BASELINE)
Table 1 and Figure 1 show the costs and energy benefits of HEC compared to traditional cloud-based solutions.
@@
According to the numerical analysis, the centralized cloud energy consumption is about: 1,927 kWh/device/year. With HEC and assuming an 80 edge split, the energy consumption is approximately 674 kWh/device/year. The resulting savings with HEC is approximately: 65.
@@
With HEC and assuming an 80 edge split, the total costs of hosting, bandwidth and software licenses is 66 per device per year. The resulting savings with HEC is about 200 per device per year or approximately 75.
Saving Comparison with Agentic Workloads:
For agentic workloads the percentage costs savings is the same and the energy saving is slightly less than traditional workloads (about 62) but the savings per device per year are almost an order of magnitude larger providing up to 10,000 kWh per year in energy and 1,500 in bandwidth and hosting cost per year per device which emphasizes why HEC is indispensable to help make agentic economy sustainable.